OpenStudy (anonymous):

how to solve ? lim [ (x^5-1)/(x^3-1)] x→a

OpenStudy (anonymous):

you need to find x

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

@robz8

OpenStudy (fifciol):

\[\lim_{x \rightarrow a}\frac{ x^5-1 }{x^3-1 }=\]\[\frac{ a^5-1 }{a^3-1 }\]